# Pruning Large Language Models with Semi-Structural Adaptive Sparse Training

### v2

@@ Abstract
The tremendous success of Large Language Models (LLMs) across various complex tasks relies heavily on their substantial scale, which raises challenges during model deployment due to their large memory consumption. Recently, numerous studies have attempted to compress LLMs using one-shot pruning methods. However, these methods often experience considerable performance degradation on complex language understanding tasks, calling into question the feasibility of pruning in LLMs. To address this issue, we propose a pruning pipeline for semi-structured sparse models via retraining, termed Adaptive Sparse Trainer (AST). Unlike previous one-shot pruning methods, AST incrementally transforms dense models into sparse ones by applying decay to masked weights while allowing the model to adaptively select masks throughout the training process. Furthermore, we observe that using distillation with a dense model as the teacher can prevent the sparse model from falling into local optima and accelerate convergence. In addition, we incorporate extra well-initialized parameters to further enhance model performance with minimal increase in memory footprint. AST can significantly enhance model performance, approaching the level of dense models. When applied to the LLaMA2-7B model, AST reduces the zero-shot accuracy gap between dense and semi-structured sparse models to 1.12% across multiple zero-shot tasks, utilizing less than 0.4% of the pretraining tokens. Our work demonstrates the feasibility of deploying semi-structured sparse large language models and introduces a novel method for achieving highly compressed models when combined with existing quantization techniques.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) are equipped to handle complex tasks (Devlin et al. 2018; Brown et al. 2020; Achiam et al. 2023) and exhibit emergent abilities (Wei et al. 2022) due to their expanding parameter count. However, this expansion poses challenges to the deployment of current state-of-the-art models, as inference speed is constrained by increasing memory latency, leading to significant computational and memory requirements. This has spurred a series of efforts to develop effective model compression techniques to reduce memory footprint and lower deployment barriers.
@@
Model pruning (Frantar and Alistarh 2023; Han, Mao, and Dally 2016; Sun et al. 2023) is an effective technique for compressing LLMs by setting a proportion of weights to zero, thereby adhering to a specific sparsity pattern. Recently, N:M sparsity has emerged as a type of semi-structured sparsity that offers an optimal balance between precision and hardware efficiency. Specifically, N:M sparsity retains only N nonzero elements out of every group of M elements. This sparsity pattern can accelerate both matrix multiplication and memory access, potentially enhancing the performance of both pre-filling and decoding processes on off-the-shelf GPUs.
Despite the promise of N:M pruning, current state-of-the-art methods for LLMs, such as SparseGPT (Frantar and Alistarh 2023) and Wanda (Sun et al. 2023), employ a post-training approach that determines the sparsity pattern layer by layer without requiring backpropagation. Although this approach improves efficiency, it often leads to significant performance degradation, particularly in knowledge-intensive tasks (Nowak et al. 2024), raising concerns about the feasibility of pruning LLMs. Additionally, while retraining pruned models has been successful in the pre-LLM era (Wang, Wohlwend, and Lei 2019; Lee, Ajanthan, and Torr 2019; Evci et al. 2020), its application to models with billions of parameters remains underexplored.
In spite of its potential, retraining sparse LLMs presents several unique challenges: (1) Retraining is computationally expensive, particularly when compared to training-free pruning methods, necessitating techniques that ensure rapid convergence. (2) The absence of original pretraining datasets often forces reliance on weaker open-source datasets for retraining, which may not fully capture the knowledge encoded in the original model. (3) The model must conform to the strict requirements of semi-structured sparsity, which adds additional constraints to the retraining process. (4) The pruning of LLMs can result in the loss of crucial language understanding and reasoning abilities, which are considerably more difficult to recover than metrics like perplexity.
@@
To overcome these challenges, we propose a novel and lightweight training method, Adaptive Sparse Trainer (AST), designed to produce high-performance sparse LLMs. For the first time, we demonstrate that semi-structured sparse LLMs can achieve competitive performance not only in terms of perplexity but also in more demanding knowledge-intensive tasks, making them viable for practical deployment. AST integrates the transition from dense to sparse models directly into the training process, gradually decaying unimportant weights to zero through a carefully designed decay mechanism. This approach allows for the revival of pruned weights during training, enabling a smooth and dynamic adjustment of the model’s connectivity pattern while adhering to the N:M sparsity structure. Furthermore, AST applies knowledge distillation using a dense model as the teacher, which can significantly speed up model convergence and prevent sparse models from getting stuck in local optima. It also helps the sparse model retain the valuable world knowledge and performance characteristics of the original dense model, thereby enhancing its generalization ability and compensating for the use of weaker training datasets. To further enhance performance, we introduce the use of a small number of carefully initialized parameters. By combining the RedPajama dataset with dynamic batching (Xia et al. 2024), AST effectively maximizes the potential of sparse models. When applied to the LLaMA2-7B model, AST achieves a 2:4 sparse configuration with minimal performance loss, demonstrating that compressed models can still perform effectively in practice. Our model further benefits from AWQ quantization, achieving competitive compression rates with state-of-the-art performance.
@@
  • We propose Adaptive Sparse Trainer (AST), a novel semi-structured pruning framework designed to compress large language models efficiently while maintaining high performance.
@@
  • We introduce SLoRB, a technique that boosts model performance by adding a small number of well-initialized parameters.
  • Applied to the LLaMA2-7B model, our method achieves a 2:4 sparse configuration with only a 0.6 increase in perplexity and a 1.2% accuracy drop in zero-shot tasks, using less than 0.4% of the original pretraining tokens.
@@ 2 Related Work
Pruning is a crucial technique aimed at reducing model size and computational costs. It originated from methods like Optimal Brain Damage (LeCun, Denker, and Solla 1990) and Optimal Brain Surgeon (Hassibi, Stork, and Wolff 1993). Based on sparsity patterns, pruning methods can be broadly categorized into unstructured, structured, and semi-structured pruning. Unstructured pruning removes individual weights (Han, Mao, and Dally 2016; Paul et al. 2023), which can maintain performance even with high sparsity. However due to its random pattern, unstructured models are difficult to accelerate. Structured pruning (Liu et al. 2017; Molchanov et al. 2019; Nova, Dai, and Schuurmans 2023; Shen et al. 2022), on the other hand, removes entire neurons, filters, or attention heads, resulting in models that are easier to accelerate on standard hardware but often suffer from severe performance loss. Semi-structured sparsity (e.g., N:M sparsity) (Hubara et al. 2021) has been applied as a trade-off between performance and achieving actual speedup. Recently a series of works (Frantar and Alistarh 2023; Sun et al. 2023; Zhang et al. 2024a, b) have made progress in pruning LLMs with billions of parameters. However, these pruned models from training-free methods still fall short in complex zero-shot performance.
@@
Another line of research (Han et al. 2015; Singh and Alistarh 2020; Renda, Frankle, and Carbin 2020; Zhou et al. 2023) has focused on retraining pruned models to enhance performance. While retraining has been shown to be effective for smaller models and simpler tasks (Kurtic and Alistarh 2022; Zhu and Gupta 2018), it often involves additional training steps (Frankle and Carbin 2019) or introduces extra parameters (Shi et al. 2023) for pruning, which limits its application to large-scale models due to the substantial computational resources required. Other retraining methods focus on unstructured sparsity, which may struggle when given stricter sparsity patterns (Lee, Ajanthan, and Torr 2019; Evci et al. 2020). Recently, Sheared LLaMA (Xia et al. 2024) employed a two-stage structured pruning process to train models that outperform others of similar size. However, they employ a computationally intensive pruning stage unsuitable for more fine-grained sparsity. Our work proposes a lightweight pruning process that enables the retraining of semi-structured sparse large language models with minimal training costs.
@@ 3 Methods
We begin by revisiting previous approaches to training sparse models, before introducing our method, the Adaptive Sparse Trainer (AST), which is described in detail. The full training pipeline is illustrated in Figure 1.
3.1 Sparsity Training
@@
where $m(\cdot)$ is a mapping that selects a mask based on the weight matrix $W$. In this work, we focus on N:M sparsity (Hubara et al. 2021), where the sparsity constraint ensures that there are $N$ nonzero elements in every $M$ consecutive weights connected to the same output.
However, when implementing the backward pass for a sparse model with automatic differentiation, the gradient cannot flow through the discrete mask $m(W)$. Previous work addresses this issue by using the straight-through estimator (STE) (Bengio, Léonard, and Courville 2013), which allows the gradient to pass through to the weights $W_{t}$ by ignoring the masking operation in the backward pass. This updates the parameters using the gradient with respect to the masked weight $\tilde{W_{t}}$, where $W_{t}$ is the parameter at iteration $t$:
@@
One common approach for retraining LLMs with specific sparsity patterns, utilizing the Straight-Through Estimator (STE) (Sun et al. 2023), involves freezing the sparsity mask at the start of training and updating only the unmasked weights using the pretraining loss function, as depicted in Figure 1. However, we identified two significant limitations with this method. First, it discards the valuable information embedded in the pruned weights and prevents a smooth transition from a dense to a sparse model, resulting in slower convergence and suboptimal performance. Second, relying solely on the language modeling loss during retraining increases the risk of the sparse model becoming trapped in a local optimum, yielding poor outcomes.
@@ 3.2 Adaptive Mask Learning with Annealing SR-STE
As highlighted in previous work (Frankle and Carbin 2019; Evci et al. 2020), the connectivity pattern is as crucial as the weights themselves in determining the performance of sparse models. Therefore, we enable the model to adaptively learn the optimal mask during training instead of fixing it permanently. Specifically, we recalculate $m(W_{t})$ based on the magnitude criteria every $\Delta t$ iterations and add a decay term to the masked weights. The intuition behind this approach is straightforward: important weights typically receive strong gradients of the same sign during training, causing their magnitude to grow despite the decay, and they will eventually be revived. In contrast, less important weights tend to receive weaker, mixed gradient signals, leading them to decay to zero and remain masked under our settings.
@@
where $\lambda_{W}(t)$ denotes the decay factor, $\overline{m(W_{t})}$ denotes the mask for pruned weights at iteration $t$, $T_{1}$ denotes the total number of training batches, and $T_{0}$ represents the batch at which the decay factor stops increasing. Compared with STE in Equation 1, masked weights receive a decay signal proportional to the decay factor $\lambda_{W}(t)$, which starts small at the beginning and increases later on.
This method, termed Annealing SR-STE (ARS-STE), is adapted from the standard SR-STE (Zhou et al. 2021). While their work focuses on pretraining settings and is primarily used to prevent mask oscillation, we apply it in the retraining context to filter out important weights. Applying ARS-STE has been shown to improve model performance compared to naive SR-STE. We also provide a detailed analysis of mask oscillation behavior during training in section 6.7.
We also experimented with different one-shot pruning criteria that incorporate activation information (Frantar and Alistarh 2023; Zhang et al. 2024a) for mask selection. However, we observed that the classic magnitude criterion consistently outperforms these methods during training, both in terms of computational efficiency and performance. Activation-based methods involve higher computational costs compared to magnitude-based masks. Moreover, their reliability diminishes as fluctuations in activation levels due to weight updates undermine their effectiveness, particularly in the later stages of training.
Given that training LLMs typically involves the use of the AdamW optimizer, we apply decay directly to the weights, as demonstrated in recent work (Hu et al. 2024). The detailed expressions and explanations are also included in Section 6.6.
@@ 3.3 Alleviating the Retraining Dilemma through Knowledge Distillation
We observe that as Retraining Dilemma is more pronounced in smaller models; therefore, we typically apply a larger $\alpha$ for these models. We also experimented with various distillation methods that utilize intermediate information for distillation. However, we found that adding constraints to intermediate outputs hinders the model’s generalization ability and leads to undesirable outcomes.
@@ 3.4 Sparse Low-Rank Boosting
Another notable aspect of our method is the initialization strategy. Classic LoRA typically employs random initialization techniques, such as Xavier initialization (Glorot and Bengio 2010) or Kaiming initialization (He et al. 2015). However, in our approach, we leverage the pruned weights as additional information to initialize the adapter weights, thereby accelerating the training process. Given that in 2:4 sparsity, each neuron can only access half of the inputs, we find that incorporating additional information to retain the weight’s first-order information can help preserve the model’s capacity. Specifically, for a weight matrix $W$ of size $N\times d$ and its corresponding mask matrix $M$, we select the rank $r$ as $\frac{d}{k}$, where $k$ can be 64, 32, 16, and so on. A smaller $k$ improves performance but also increases memory usage. We select a projection matrix $X$ of size $\frac{d}{k}\times d$ and a SLoRB weight matrix $S$ of size $N\times\frac{d}{k}$. The matrices $X$ and $S$ are defined as follows:
@@
We define Group $G_{ij}$ as the elements from $j\cdot k$ to $(i+1)\cdot k-1$ in row $i$. As illustrated in Fig 3, each SLoRB weight $S_{ij}$ is broadcast within Group $G_{ij}$. By setting $S_{ij}$ as the mean of all pruned weights in Group $G_{ij}$, this design ensures that the total mean of groups $G_{ij}$ remains consistent after pruning. Our initialization method provides the model with a better starting point in training. Although different initialization methods may ultimately converge to similar outcomes given sufficient training steps, our method converges more rapidly, making it particularly advantageous when the computational budget is limited, as demonstrated in Section 6.4. We refer to this method as Sparse Low-Rank Boosting (SLoRB), which trades off memory overhead for enhanced performance. The complete pseudocode for our method is provided in Algorithm 1.

### v3

@@ Abstract
The remarkable success of Large Language Models (LLMs) relies heavily on their substantial scale, which poses significant challenges during model deployment in terms of latency and memory consumption. Recently, numerous studies have attempted to compress LLMs using one-shot pruning methods. However, these methods often suffer from considerable performance degradation on complex language understanding tasks, raising concerns about the feasibility of pruning in LLMs. To address this issue, we propose Adaptive Sparse Trainer (AST), a novel and efficient retraining framework tailored for semi-structured sparse models. AST enables models to learn optimal masks during the weight update process without incurring additional computational overhead. Furthermore, we demonstrate that incorporating knowledge distillation significantly improves retraining efficiency and enhances model performance under fixed computational constraints. Additionally, a supplementary set of well-initialized parameters is integrated to further augment the model’s efficacy. AST achieves state-of-the-art performance with minimal training cost. When applied to the LLaMA2-7B model, AST reduces the perplexity and zero-shot accuracy gap between dense and 2:4 semi-structured sparse models to 0.6 and 1.16%, respectively, utilizing less than 0.4% of the pretraining tokens and GPU hours. Our work demonstrates the feasibility of deploying semi-structured sparse LLMs and offers a promising alternative for achieving highly compressed models when combined with existing quantization techniques.
Codehttps://github.com/thu-ml/Adaptive-Sparse-Trainer
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) are equipped to handle complex tasks (Devlin et al. 2018; Brown et al. 2020; Achiam et al. 2023) and exhibit emergent abilities (Wei et al. 2022) due to their expanding parameter count. However, this continual growth in model size poses significant challenges for practical deployment. In particular, inference speed suffers due to the increasing computational and memory demands. This has spurred a series of efforts to develop effective model compression techniques aimed at reducing memory footprints and easing the constraints associated with deploying these large-scale models.
@@
Model pruning (Frantar and Alistarh 2023; Han, Mao, and Dally 2016; Sun et al. 2023) is an effective technique for compressing LLMs by setting a proportion of weights to zero, thereby adhering to a specific sparsity pattern. Recently, N:M sparsity has emerged as a type of semi-structured sparsity pattern that offers an optimal balance between precision and hardware efficiency. Specifically, N:M sparsity retains only N nonzero elements out of every group of M elements. This sparsity pattern can accelerate both matrix multiplication and memory access, potentially enhancing the performance of both pre-filling and decoding processes on off-the-shelf GPUs. Despite the promise of N:M sparsity, current state-of-the-art methods for pruning LLMs, such as SparseGPT (Frantar and Alistarh 2023) and Wanda (Sun et al. 2023), employ a post-training approach that determines the sparsity pattern in a layer-by-layer fashion without back-propagation. Although these methods improve efficiency, they can lead to significant performance degradation, particularly in knowledge-intensive tasks (Nowak et al. 2024), raising concerns about the feasibility of pruning LLMs. Additionally, while retraining pruned models has been successful in the pre-LLM era (Wang, Wohlwend, and Lei 2019; Lee, Ajanthan, and Torr 2019; Evci et al. 2020), its application to models with billions of parameters remains under-explored.
@@
Although retraining sparse LLMs holds significant potential, it introduces several unique challenges: (1) Retraining is computationally expensive, necessitating techniques that ensure rapid convergence. (2) The output model must adhere to strict sparsity pattern, which adds additional constraints to the retraining process. (3) In order to achieve optimal performance, both masks and weights should be learnable during training. (4) Pruning LLMs may compromise essential language understanding and reasoning abilities, which are significantly harder to restore compared to simpler metrics such as perplexity. To this end, we propose a novel and efficient training method, Adaptive Sparse Trainer (AST), designed to produce high-performance sparse LLMs. For the first time, we demonstrate that pretrained 2:4 sparse LLMs can achieve competitive performance not only in terms of perplexity but also in more demanding knowledge-intensive tasks, making them viable for practical deployment. AST integrates the transition from dense to sparse models directly into the training process, gradually decaying unimportant weights to zero through a carefully designed decay mechanism. This approach allows for the revival of pruned weights during training, enabling a smooth and dynamic adjustment of the model’s connectivity pattern while adhering to the N:M sparsity structure. Furthermore, AST applies knowledge distillation using dense model as the teacher, which can significantly speed up model convergence and prevent sparse models from settling into local optima. It also helps the sparse model retain the valuable world knowledge and performance characteristics of the original dense model, thereby enhancing its generalization ability and compensating for the use of weaker training datasets. To further enhance performance, a supplementary set of parameters is integrated using information from pruned weights. When applied to the LLaMA2-7B model, AST achieves a 2:4 sparse configuration with minimal performance loss, demonstrating that compressed models can still perform effectively in practice. Our model further benefits from AWQ quantization, achieving competitive compression rates with state-of-the-art performance.
@@
  • We propose Adaptive Sparse Trainer (AST), an novel semi-structured pruning framework designed to compress large language models efficiently while maintaining high performance.
@@
  • We introduce SLoRB, a technique that boosts model performance by adding a supplementary set of well-initialized parameters.
  • When applied to LLaMA2-7B model, our 2:4 sparse model experiences only a 0.6 increase in perplexity and a 1.16% accuracy drop in zero-shot tasks, with less than 0.4% of the pretraining cost.
@@ 2 Related Work
Pruning is a prevalent technique aimed at reducing model size and computational costs. It originated from methods like OBD (LeCun, Denker, and Solla 1990) and OBS (Hassibi, Stork, and Wolff 1993). Based on sparsity patterns, pruning methods can be broadly categorized into unstructured, structured, and semi-structured pruning. Unstructured pruning removes individual weights (Han, Mao, and Dally 2016; Paul et al. 2023), which can maintain performance even with high sparsity. However, due to its random pattern, unstructured models are difficult to accelerate. Structured pruning (Liu et al. 2017; Molchanov et al. 2019; Nova, Dai, and Schuurmans 2023; Shen et al. 2022), on the other hand, removes entire neurons, filters, or attention heads, resulting in models that are easier to accelerate on standard hardware but often suffer from severe performance loss. Semi-structured sparsity (e.g., N:M sparsity) (Hubara et al. 2021) has been applied as a trade-off between performance and achieving actual speedup. Recently, a series of works (Frantar and Alistarh 2023; Sun et al. 2023; Zhang et al. 2024a, b) have made progress in pruning LLMs with billions of parameters. However, these pruned models from training-free methods still fall short in complex zero-shot performance.
@@
Another line of research (Han et al. 2015; Singh and Alistarh 2020; Renda, Frankle, and Carbin 2020; Zhou et al. 2023) has focused on retraining pruned models to enhance performance. While retraining has been shown to be effective with smaller models and simpler tasks (Kurtic and Alistarh 2022; Zhu and Gupta 2018), it often involves additional training steps (Frankle and Carbin 2019) or introduces extra parameters for the pruning process (Shi et al. 2023), which limits its application to large-scale models due to the substantial computational resources required. Other retraining methods focus on unstructured sparsity, which may struggle when given stricter sparsity patterns (Lee, Ajanthan, and Torr 2019; Evci et al. 2020). Recently, Sheared LLaMA (Xia et al. 2024) employed a two-stage structured pruning process to train models that outperform others of similar size. However, they employ a computationally intensive pruning stage, unsuitable for more fine-grained sparsity. Our work proposes a lightweight pruning process that enables the retraining of semi-structured sparse large language models with minimal training costs.
@@ 3 Methods
We begin by revisiting a naive approach to training sparse models before introducing our method. The full training pipeline is illustrated in Figure 1.
3.1 Sparse Training
@@
where $m(\cdot)$ is a mapping that selects a mask based on the weight matrix $W$. In this work, we focus on N:M sparsity (Hubara et al. 2021), where there are $N$ nonzero elements in every $M$ consecutive weights in the same row. However, when implementing the backward pass for a sparse model with automatic differentiation, the gradient cannot flow through the discrete mask $m(W)$. A straightforward approach to solve this issue is the straight-through estimator (STE) (Bengio, Léonard, and Courville 2013) which updates the parameters using the gradient with respect to the masked weight $\tilde{W_{t}}$, where $W_{t}$ is the parameter at iteration $t$:
@@
A previous method (Sun et al. 2023) employs a fixed mask derived from one-shot pruning techniques and updates only the remaining parameters using language modeling loss. However, this strategy has two significant limitations. First, it discards the valuable information embedded in the pruned weights and prevents a smooth transition from a dense to a sparse model, resulting in slower convergence and suboptimal performance. Second, relying solely on the language modeling loss during retraining increases the risk of the sparse model becoming trapped in a local optimum, leading to poor outcomes.
@@ 3.2 Adaptive Mask Learning with Annealing SR-STE
As highlighted in previous work (Frankle and Carbin 2019; Evci et al. 2020), the connectivity pattern is as crucial as the weights themselves in determining the performance of sparse models. Therefore, we enable the model to adaptively learn the optimal mask during training instead of fixing it permanently. Specifically, we recalculate $m(W_{t})$ based on the magnitude criterion every $\Delta t$ iterations and add a decay term to the masked weights. The intuition behind this approach is straightforward: important weights typically receive strong gradients of the same sign during training, causing their magnitude to grow despite the decay, and they will eventually be revived. In contrast, less important weights tend to receive weaker, mixed gradient signals, leading them to decay to zero and remain masked under our settings. We build upon SR-STE (Zhou et al. 2021) by applying L2 decay to masked weights, as this approach helps preserve overall model performance.
@@
where $\lambda_{W}(t)$ denotes the decay factor, $\overline{m(W_{t})}$ denotes the mask for pruned weights at iteration $t$, $T_{1}$ denotes the total number of training batches, and $T_{0}$ represents the batch at which the decay factor stops increasing. Compared with STE in Equation 1, masked weights receive a decay signal proportional to the decay factor $\lambda_{W}(t)$, which starts small at the beginning and increases later on. This method, termed Annealing SR-STE (ARS-STE), has been shown to improve model performance compared to naive SR-STE. We also provide a detailed analysis of mask oscillation behavior during training in Section 7 of the Appendix.
We experimented with various one-shot pruning criteria incorporating activation information (Frantar and Alistarh 2023; Zhang et al. 2024a) for mask selection. However, the classic magnitude criterion consistently outperformed these methods in both computational efficiency and performance. As activation-based approaches incur higher computational costs and are less reliable due to activation fluctuations caused by weight updates, leading to degraded performance in later training stages.
Since the naive SR-STE is only compatible with the SGD optimizer, we introduce further adjustments to support more advanced optimizers such as AdamW. Specifically, as AdamW maintains both first-order and second-order moments of the gradient, we decouple the weight decay term from the first-order moment instead of directly adding decay to the gradient, as done in (Hu et al. 2024). This separation ensures that the weight decay term depends solely on the current weight, avoiding interference with momentum calculations. We then use the decoupled first-order signal to update the second-order moment. Detailed expressions and explanations are provided in Section 6 of the Appendix.
@@ 3.3 Alleviating the Retraining Dilemma through Knowledge Distillation
We observe that the Retraining Dilemma is more pronounced in smaller models; therefore, we typically apply a larger $\alpha$ for these models. We also experimented with various distillation methods that utilize intermediate information for distillation. However, we found that adding constraints to intermediate outputs hinders the model’s generalization ability and leads to undesirable outcomes.
@@ 3.4 Sparse Low-Rank Boosting
Another notable aspect of our method is the initialization strategy. Classic LoRA typically employs random initialization techniques, such as Xavier initialization (Glorot and Bengio 2010) or Kaiming initialization (He et al. 2015). However, in our approach, we leverage the pruned weights as additional information to initialize the adapter weights, thereby accelerating the training process. Given that in 2:4 sparsity, each neuron can only access half of the inputs, we find that incorporating additional information to retain the weight’s first-order information helps preserve the model’s capacity. Specifically, for a weight matrix $W$ of size $N\times d$ and its corresponding mask matrix $M$, we select the rank $r$ as $\frac{d}{k}$, where $k$ can be 64, 32, 16, and so on. A smaller $k$ improves performance but also increases memory usage. We select a projection matrix $X$ of size $\frac{d}{k}\times d$ and a SLoRB weight matrix $S$ of size $N\times\frac{d}{k}$. The matrices $X$ and $S$ are defined as follows:
@@
We define Group $G_{ij}$ as the elements from $j\cdot k$ to $(i+1)\cdot k-1$ in row $i$. As illustrated in Fig 3, each SLoRB weight $S_{ij}$ is broadcast within Group $G_{ij}$. By setting $S_{ij}$ as the mean of all pruned weights in Group $G_{ij}$, this design ensures that the total mean of groups $G_{ij}$ remains consistent after pruning. Although different initialization methods may ultimately converge to similar outcomes given sufficient training steps, our method converges more rapidly, making it particularly advantageous when the computational budget is limited, as demonstrated in Section 4 of the Appendix. We refer to this method as Sparse Low-Rank Boosting (SLoRB), which is an optional technique that trades off memory overhead for enhanced performance. The complete pseudo-code for our method is provided in Algorithm 1.